OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. d = 13 A = 265.33 sq. units correct me if im wrong?..

OpenStudy (anonymous):

d=13 therefore \[r= \frac{13}{2}\] \[Area-of-circle= \pi r^2 = 3.14\times (\frac{13}{2})^2=3.14\times \frac{169}{4}\] \[=\frac{530.66}{4}=132.665 units^2\] Youer ans doesn't matches with this one > i think that you have substituted d=13 which is wwrong that is diameter convert it into radius dividing it by 2 i.e. \[r= \frac{d}{2}\]

OpenStudy (anonymous):

dam im really not good at these.. :/ thank you